Today I tuned the mode matching into the south cavity by adjusting the two lens translation stages. The 64.4-mm ROC lens is at 15.63 mm, and the 38.6-mm ROC lens is at 4.01 mm. We currently have 80% visibility (see first attachment). One of the lenses is at the end of the range of its stage, so we will have to work to reposition it and continue mode matching.
Tara was able to tame the cavity servo loop. The third attachment is the result of several SR785 measurements of the error signal power spectrum. I converted this to a frequency noise power spectrum (fourth attachment) by extracting the voltage/frequency calibration factor from the error signal as follows.
I smoothed the PDH error signal by convolving with a gaussian and then numerically differentiated (second attachment). The locations of the carrier and the sidebands can then be read off, along with the magnitude of the derivative at zero cavity offset. Sidebands are at −14.5 ms and +14.7 ms, and the carrier is at 0.2 ms. Since the sidebands are 14.75 MHz from the carrier, this gives the time-voltage calibration as 1.01 GHz/s. The zero crossing of the carrier has derivative 16600 V/s, and hence the PDH slope is 61 kHz/V.
I dialed the 38.6-mm ROC lens back to 1.11 m, twiddled the periscope mirrors, and got only marginal improvement in the visibility.